OpenStudy (anonymous):

the foci of an ellipse are (0,+ or - 4) and the length of the major axis is 12. How long is the minor axis?

OpenStudy (anonymous):

check this one out! it will help with ellipse all together: http://home.windstream.net/okrebs/page62.html

OpenStudy (anonymous):

From the coordinates of the foci, c = 4 and the major axis is the y axis. From the length of the major axis, we obtain a = 6. Also b^2 = a^2 - c^2 =36-16= 20. The equation of the ellipse is given by x 2 / 2sqrt(5) + y 2 / 36 = 1